Hall Microscopes Facility
What's Down & Scheduled Repairs
On Going Issues:
Some users have been having trouble connecting to
\\Minsky. If you encounter this situation use this i.p. number
instead \\184.108.40.206 or \\minsky.med.unc.edu
- VPN Outage: Thursday 19 March the campus VPN will be down for
~15 minutes starting at ~6 AM. This is to upgrade the
capacity of the UNC VPN system. During this time existing
VPN connections will be lost. E.g. Remote Desktop connections will be lost, Cisco UNC-VPN
sessions will become disconnected.
02 - Calendar
server is currently down and unreachable. A support
ticked has been submitted in order to get the virtual machine
Leica SP2 Confocal - The
microscope stand is not initializing. The system is not useable.
The microscope stand is being assessed (i.e. dismantled)
- Leica SP2 Confocal - The UV AOTF cooling fan bearing is failing. Please
power this up as needed and down when finished. A new fan is being
- Sep 23: Leica SP2 - Z-galvo wires replaced. Stage should
have full range (+/-80 um) without oscillation. Stainless wire 7
filaments of 100 um each. Fraying wire filaments caused oscillations in the z-positioning
- Sept 03, ~11 AM to ~1 PM - The calendar pages were down due to a
failure of the virtual public server hardware, NYC. Sorry
for the inconvenience.
- Volocity license server is currently unavailable due to the
recent firewall SOM setup around the Cell Biology subnet
- Wednesday April 22, 10:30 AM - \\Minsky has been moved to
Marsico Hall. File shares were not accessible while the hardware
was being moved and Ethernet port activation was pending.
- Sep 24: The DNS name entry of this web site was erased. OIS
restored the entry Sep 25. The web site could still be access with
the i.p. http://220.127.116.11
during this time.
- Sep 24: The DNS name entries of several microscope computer
systems were erased due to unknown reasons. OIS restored the names
- Sep 11 & 12: This web site's server had to be
restored to its Aug 18 state, including the booking calendar
database, due to an unrecoverable crash. Calendar database backups
have been confirmed to be running again.
- Aug 16: Power outage is scheduled for 8 AM to 12
Noon, Saturday. Webserver and the file and Volocity server
should continue to run during this period since emergency power is not
scheduled to be turned off. Workstations and microscopes will not be usable
when the power is out.
- Aug 04 at ~5:30 PM - The file server \\Minsky was down.
This server also runs the Volocity
Licence (British spelling) Server. Therefore both
file transfers can not be done and Volocity could not be used. The
issue can only be checked during normal work hours.
Sorry for the delay in restoring the services.
Service was restored in the AM of Aug 05.
- June 17: Due to an unexpected glitch the power remained
out for ~2 hrs.
- June 17: Due to scheduled building power outage and
emergency power generator test most computers and backup RAID servers will
be powered down Monday evening. Users may have to manually power up
computers that run acquisition systems Tuesday morning.
- Powered down are:
- Olympus FV1000 computer - \\Sutton
- Leica SP2 computer - \\Laplace
- Zeiss 510 computer - \\Hertz
- Upright Nikon computer - \\Hooke
- Leica inverted color camera computer - \\Snell
- Leica inverted b/w camera computer - \\Lister
- Fluorescence dissecting scope - \\Gould
- Spinning disk/wide field inverted Nikon - \\Nipkow
- Powered down and of no immediate impact to users:
- Raid array #1 Raid & computer \\Maxwell
- Raid array #2 Roach & \\Ai
- Raid array #3 Ant & \\Abbe
- Backup webserver
- Web & calendar server http://microscopy.unc.edu
- Minsky file server
- May 28: File server \\Minsky temporarily lost network connectivity from
6:49 PM to 7:00 PM due to Ethernet wall jack being damaged.
- Apr 25: \\Lister - A new computer has been swapped into \\Lister's role.
All users will have to restore there user profile in order to capture images
using the Photometric HQ2 ccd camera with C-Imaging.
- Apr 21: Incubator - Humidity control is giving alarms. This is a
controlled humidity unit and it may not be possible to simply place water in
the try on not have the humidity alarm go off. Note that temperature
and CO2 control are not affected by the humidity control problem.
- Apr 19: \\Lister - Looks like the mother board has blown out (just
thought the system had the XP update bug)
- Apr 17: These workstations
have succumbed to the Windows XP update bug (KB2952678) and were off
- Fluorescence dissecting
scope (\\Gould) [fixed]
- Inverted Leica widefield
monochrome camera (\\Lister) [fixed]
- Inverted Leica widefield
color camera (\\Snell) [fixed]
- Leica SP2 confocal (\\Laplace) [fixed]
- Win 32 Image Processing Workstation \\Nomarski
- Other possible system that could be
affected: Zeiss 510 (\\Hertz), Spinning
- Apr 10: Web server &
calendar crashed. Back in service 10:40 AM - Nothing to do
- Feb 03: \\Snell, the computer for the color camera at the Leica inverted
widefield microscope, hard disk drive was replaced and restored with a
system image backup of Feb 21, 2013. The system is being updated. Please
help us by letting us know if there are anomalies, things missing on Windows
and if it does not work as it should, thanks.
- Feb 01: The MHmicroscopy AD/DC server did not restart after a brief
power outage at 7:01 AM. Logons may not have be possible at the key boards
of facility computers or over the network to servers/shared directories.
Server was manually restarted at about 4 PM, after the UPS was replaced.
- Jan 23: Incubator repair done. Conditions are once again
controlled 37C, 5% CO2 & 95% and the readings are steady. Fyrite
calibration has been carried out.
- Sep 13: Leica SP2 Confocal: System - The hard disk drive has crashed.
Due to the difficulty of obtaining a drive the whole computer will be
- Sep 13: Leica SP2 Confocal:
The system is unavailable due to a hard disk drive crash. Data
from prior to Sep 13 is available at \\minsky.med.unc.edu\LeicaSP2.
- Sep 06: Leica SP2 Confocal: System seems to be working.
galvanometer controller cards and can not reproduce problem.
Please use the system, and we will be checking with you. Of course
please let us know about issues.
Sep 04: Leica SP2 Confocal:
System is not useable due to bad x-axis galvo. Engineer
is expected Sep 06.
Sep 04: \\Minsky file server - 7:20 to 9:35 PM file server
was inaccessible due to an unplugged network cable, which had a
missing locking flange. The cable has been replaced.
Apr 19: Leica SP2 confocal - AOBS fans have been replaced.
Mar 27: Olympus FV1000 confocal - the replacement red 638 nm red laser
has been installed. All laser in the merge box were alignment. All laser
lines are functioning again.
Mar 26: The AD/DS mhmicroscopy (mh-microscopy.med.unc.edu) domain
server (Marvin) had a failed mirror drive. The system did not
reboot to the remaining mirror until the sATA cables were reversed.
The mirror pair has been rebuild.
Mar 19: Power will be turned off
briefly ~30 mins in the evening (sometime between ~9 to ~11 PM). This
is all part of standard University power distribution maintenance.
Mar 19 Laptop wireless access to \\Minsky file
shares is not working. This is due to network infrastructure issues
beyond our responsibility/control. Access from desktops and laptops plugged
into the wired network (Ethernet) works normally.
Mar 05: Olympus
FV1000 Confocal Microscope - red laser has failed ( 635 nm for far red, e.g. Alexa
633/647, Draq-5, ToPro-3). Note that the other
lasers are unaffected, so dyes such as 405 nm -> DAPI,
440 nm -> CFP,
488 nm -> Alexa488/GFP/FITC, 559 nm -> Rhodamine/TexasRed/Alexa 555/568, etc.
can still be imaged.
The web server and calendars were down from 6:20 AM to 9:55 AM. The UPS was
found off again. The UPS was replaced.
The web server and calendars were down from 7:05 AM to 8:58 AM. The UPS was
found off. We are investigating why it turned off.
Feb 04: Network disruption continue. We need your help in tracking down
problems, so let us know if you have issues connecting over the network by
telling us which compouter(s) you are trying to connect to, your location
and whether you are using a desktop or laptop.
Feb 01: Expect network weirdness.
This is due reconfiguring
of the VLAN in Thurston Bowles. Problems will roll around
different network jacks and hence the associated computers.
Once the reconfiguration is finished the weirdness factor should
almost go away completely. If you have problems please call so
that we can know that there is a problem, and we will try to find a
workaround for you.
Jan 31: On Thursday evening several
computers could not connect to several other computers in 6th floor
of Thurston Bowles building. The disruption lasted about 30
minutes, and should be taken as a warning not to scan images
directly to servers over the network. ITS/OIS was notified and
a ticket was created.
- Jan 22: Leica SP2 AOBS confocal - The computer keyboard has been
replaced. The 1, q, a
& z keys have stopped working, perhaps due to a drink drink
spill caused . Since an US English keyboard has replaced the original
German keyboard the following keys are remapped:
shift + \
shift + ‘
we're looking for the key combination
- Jan 16: Leica SP2 AOBS confocal -
UV correction lens for imaging DAPI with the 63x 1.4 NA
now switches into place automatically. Laser UV correction
lenses were aligned. Will need to determine if field illumination is
adequate at 10x,
20x, 40, & 63x. Please let us know if you notice that the UV field
illumination is not even.
- Jan 11: Leica SP2 AOBS confocal - Optical mouse failed. It was
- Sep 23: Olympus FV1000 Confocal: The C: drive on
\\Sutton failed completely. The system has been restored to the May 24,
2011 state. User short cuts to \\Minsky, the users directory, default
web page, etc. will need to be reestablished for each user. See
Michael or Neal for assistance. Note user data should be intact on the
- Sep 19: \\Nomarski Image
Processing Working station has suffered an irrecoverable hard disk
drive failure. The system is not usable and is currently only
restored to its May 2011 state. Software will be reinstalled
over the next week or so. Let Michael know ASAP if you need a
particular piece of software sooner.
- Jul 28: Leica SP2 AOBS confocal - The engineer is expected at the end of
this week. The 488 nm laser will be addressed and the sticking UV correction
lens for imaging DAPI with the 63x 1.4 NA objective. The UV correction lens
for 40x 1.25 NA oil and 10x/20x objectives do switch in correctly.
July 12: There is an apparent after hours power failure in
the server room. It will not be possible to logon to
MHmicroscopy computers including \\Minsky and other file
shares. The duration of the power loss has exceeded the run
time of the UPS batteries. An on campus visit is
required to assess the situation.
- Jun 28: Leica SP2 Confocal
Microscope is running. Note: User profiles may be different
and network access is restricted to \\Minsky
only and will need to be setup when users login, or at your request
Michael or Neal can setup this up and issue a new password.
- Jun 23:
This Web Server &
associated booking calendars will be
run off the backup server. Page updates may be slow.
Contact Michael (843-3268 anytime) if you have booking or other
- Jun 18: Leica SP2 confocal
microscope. The C: drive has failed. We are expecting
delivery of a replacement drive on Wednesday Jun 20. When this drive
arrives its operational worthiness needs to assessed before the WindowsNT4
operating system, drivers and confocal software can be reinstalled. The
computer/microscope is currently unusable.
Michael to access data up to 3 AM June 12, which is stored on the
- Jun 12: Leica SP2 confocal
microscope: The C: drive has failed. We expect the
computer/microscope to be unusable for a few days. Contact
Michael in order to arrange to access data save on
\\Laplace up to 3 AM June 12, which is stored on the
- Jun 08: Olympus FV1000 MPE confocal - The system is back in
operation. A failing hard disk drive (C:) was causing the computer to
not boot into Windows. The drive was replaced, however the system was restored
to a slightly earlier version (3.00a) of FV-10 confocal software. Version 3.01 was
running with the failed drive. There is probably little, if any, consequence
using the earlier version.
Data on the D: drive is intact, therefore saved images and user scanning
profiles are the latest versions. We will need to get a copy of the
software distribution disk from Olympus in order to restore version 3.01 of
the FV-10 software. Please let Michael or Neal know if you encounter
problems with the system.
To connect to this computer over the campus network please
use this address (unc):
Just connecting to \\sutton\users will fail.
- Jun 07: Olympus FV1000 MPE confocal - The system is not
operational since the computer has failed. We are trying to determine
what the bad component is. Data files stored on the D: drive in the
last 4 days have been copied to \\minsky\users2\sutton so that they can be
- Apr 12: Multiphoton Ti:Sapphire laser is back in operation on the
Olympus FV1000 con focal microscope. The chiller unit was
replaced. Drier cartridge was replaced in March.
- Mar 19: Leica SP2 Confocal microscope - Microscope stand LCD was
showing ?? - ? for objective. LCS software was not initializing stand.
Resetting microscope stand control box did not help. Problem was solved by
reseating objective turret. Tightening screw on left of stand was not
screwed in firmly at all. It may have been loose for the life of the
scope! Note need to reprogram objective parfocality.
- Mar 22 Multiphoton Ti:Sapphire laser is down due
to a flakey chiller unit. Discussion with service is in
progress and a replacement is pending. The standard lasers are still usable on the Olympus
FV1000. Please do not turn on the Mai Tai Ti:Sapph laser.
- Feb 27: Leica SP2 Confocal Microscope - Scan head backplane and X3
connector cable replaced. This should eliminate the lower voltage
problem that has been plaguing the system for quite a while.
- Feb 23: New server firewall settings are being
implemented in order to meet standards for servers set for the Sch
Of Med. by ITS.
Problems may occur during this shakedown
period. e.g. With connections to \\Minsky files shares & shares to
other computers. e.g. Inability to log onto facility computers.
Please let Michael ASAP if you encounter problems - voice 843-3268,
text 919-912-9380, email see in Contacting
- Feb 22: Volocity License Sever - UNC initiated
security scans are stopping this service from running on the server
computer. If it is stopped
this web page will
not respond & Volocity will not run authorize on your computer. We
are manually restarting the service. Please contact Michael if
you find it unavailable so that it can be started.
- Feb 20: OIS Ethernet connections - We are soon to be knackered. See
NAC & guide
re. new SOM network requirements.
- Jan 13: Active Directory Domain Server (\\Marvin the password
computer) has been crashing for mysterious reasons. If you can not log
in or connect over the network please
call or text Michael
anytime. This computer will be upgraded.
- Jan 13: LeicaSP2 - UV Illumination is better centered.
Fiber coupling has been improved several fold. A new UV fiber is being
ordered and is expected to provide more even illumination.
- Jan 03: LeicaSP2 - The computer power supply was
replaced after damage as a result of a crash. A CPU cooler fan
is on order.
- Dec 28: Domain Controller Computer (\\MARVIN) - Power supply has failed. Log ins and shared
directories may not have been available from
~12:00 PM to ~3:30 PM. The power supply was replaced.
- Dec 15: Leica SP2 Confocal - The UV laser is available.
Please start at tube current at zero and run at no more than two
turns. Set the 364 nm AOTF to about 15% to 20%.
- Oct 11: Leica SP2: UV
laser has failed - We are searching for a replacement unit. The visible lasers are still operational and can be used
normally. Please consider using the Olympus FV1000
confocal while UV laser is being repaired. Please contact
the usual people
- Sep 19: Olympus FV1000 MPE confocal:
Motorized x-y stage has been
- Sep 19: Inverted Leica widefield scope: Computer
\\Lister - access to shared folders (\\lister\users, \\lister\users2)
is not working properly. This problem has been fixed by adding the
IRPStackSize parameter to the lanmanager service.
- Sep 16: Olympus FV1000 confocal: Stage has been repaired.
There was a broken wire at the stage connector. Laser alignment was
- Sep 16: Leica Stereo/macro-scope: - USERS1 & USERS2 hard disk
drive failed on the \\Gould computer. Failed computer power supply has
been replaced. Power supply failed after hard disk drive failed.
- Sep 16: Leica SP2: 63X
1.4 NA PL APO blue oil immersion objective is damaged after being crashed
into the condenser due to uncontrollable stage movement as a result of a MIC
controller crash. Still out for repair.
stop gap while the 63x is out for repair please use 40x and zoom in an extra
1.575x. There will be a slight (11%) loss of x-y resolution which is
not subjectively noticeable, and would only be detectable under the most
optimum of sample and imaging conditions. Note: If you will be
doing x-z scanning with UV please see Michael or Neal about an alternate 40x
1.25 NA objective which has good x-z UV colocalization.
- Sep 06: Leica SP2: UV laser is working. Used 4D Enterprise
(0) power supply.
- Jul 21: Laser Micro Dissection: Pulsed UV laser has failed.
- Aug 02: Leica SP2 confocal microscope: Service with Jim & Jim.
Lasers, voltages, etc. are OK. Perhaps AOTF is misbehaving. If
lasers cut try changing settings on AOTF controller to see if laser output
- June 03: \\Marvin the MHmicroscopy AD/domain controller
had a bad mirror disk which was causing this computer to crash.
Logging on to microscope computers and network access may have been
difficult on Friday, June 3. The defective redundant dive has been
replaced - no data has been lost.
- May 27: Leica DMIRB inverted widefield
microscope: A Hamamatsu Orca-ER CCD b/w camera has been put in
service in place of the broken Cool Snap HQ2 camera. The acquisition
software will need to be updated in order to use the monochrome camera.
This is quick to do. Please see Neal or Michael for assistance
carrying this out with your log in on the \\Lister
- May 18: Image processing workstation
hard disk drive has failed. This computer will be unavailable while
the system is being rebuilt.
- May 17: Nikon Nipkow Spinning Disk/Widefield
microscope: The motorized dichroic
filter cube turret has failed. Inspection the unit has revealed that
salty fluid entry into the unit has caused corrosion of a crucial ball
bearing. The turret rotation is stiff, rough and can not be driven by
- May 16: Leica DMIRB inverted widefield
microscope: The Cool Snap HQ2 CCD camera has become flaky.
Temporarily an EMCCD camera has been put in its place. Please run this
camera at multiplier gain 0 first then increase exposure time, &/or gain to
2 or 3. Multiplier gain will dramatically increase at over 17.
The EMCCD camera has a dramatically different overload characteristics
compared to conventional CCD cameras. Also the EMCCD only has a 512 by
512 sensor rather than the 1024 by 1024 sensor of the Cool Snap HQ2.
Calibration images of the reticule slide need to be taken with the Cascade
512 for each objective on the Leica DMIRB. When we get a chance a
Hamamatsu Orca ER will put on the system while we search for funds to repair
the Cool Snap HQ2.
- Feb 22 - Incubator 37C, 5% CO2, 95% Humidity: Is
- Feb 15 - Incubator 37C, 5% CO2,
95% Humidity: A broken plastic fitting prevents
active humidification. We are seeking a replacement
part, which may take a few days.
- Feb 11: Leica SP2 Confocal
Microscope - The UV laser has been aligned and the system is
useable. Please make sure the tube current control is not turned more than 2
turns from the minimum position.
- Feb 09: Leica SP2 Confocal
Microscope - Backup UV laser is working! We still need to get the UV laser aligned to the scope,
which will take a few days. Users of this system will receive
an e-mail regarding its status
- Feb 08: Leica SP2 Confocal microscope - The
UV laser head has failed. All other laser lines (488, 514, 561, 633
nm) are still useable. Expected replacement sometime the week of Feb
14. We have a spare laser to swap in. If this laser
functions it will need to be aligned with the fiber coupling.
- Jan 06: Zeiss 510 Meta Confocal
is available for use.
- Please double click "New User Generic
Profile" on the desktop before starting the Zeiss software for
the first/next time.
- Old data is secured on backup
servers. Restored upon request.
- For help with restoring scan configurations
or restoring old data please contact Michael or Neal
- Jan 05: Zeiss 510 Meta Confocal - Is not available due
to a computer/disk crash. We are working to get it operational. Data should be
available on daily backups. Contact Michael or Neal for
- Nov 09: Zeiss 510 - Ch3 filter wheel serviced.
- Nov 04: Zeiss 510 - Ch3 filter wheel has been replaced. Arc
lamp field stop and aperture irises will be replaced.
- Oct 11: Zeiss 510 - PMT power supply has been repaired. The
system is running normally.
- Oct 08: Zeiss 510 - PMT Ch-3 (red, far red
emission) has failed. Repair is schedule for Oct 11 morning.
See Neal or Michael about using the Ch-S (meta) as a work around.
- June 24: Leica SP2 confocal - System has been restored.
Replaced detector board, power module and adjust phase.
- June 22: Leica SP2 confocal
- System is crashing frequently. Service is scheduled for
- June 14: Leica SP2 confocal - System was serviced. New scan
head controller (SCSI?), exchanged MB but returned to original. UV
lens motor not exchanged. Scheduled for exchange driver board & main
- June 03: Leica SP2 confocal - Red laser (633 nm) has been
replaced along with its power supply. The system is running again.
- June 01: Leica SP2 confocal - Red laser
(for excitation of CY5, Alexa 633/647, Draq5,
etc.) fuse has blown again,
reiterating the indication of a bad red HeNe laser. Replacement parts
have arrived and will be installed June 2.
- May 28: Leica SP2 confocal - Red laser operation has been
restored for now. Although it can not certain how long it will remain
functioning. Please use it while you can. If it fails please let
us know by e-mail/phone. Replacement part is expected June 1 and will
be installed June 2.
- May 27: Leica SP2 confocal - Red 633 nm laser is not working.
Fuse is blown indicating a bad laser &/or power supply.
Replacement parts are expected June 2.
- April 16: Nikon
TE2000 - The computer called \\Morpheus has failed
and has been replaced with another computer called \\Ploem.
Independent of the motherboard failure the disk drive has failed. No
data was lost, and please contact Neal if you need to access any of the
backed up data files. For now, data files should be stored to
C:\USERS. Your may create a named directory for yourself. \\Ploem\users1
is the network share name of the c:\users directory.
- Important: The filter and camera configurations have been changed
slightly in the rebuilt system. Please contact Neal for help setting
up filter and camera configurations.
- April 07: Leica SP2
confocal service maintenance completed. Please log any issues.
- Feb 27: Leica SP2 confocal - The "SCSI
error" system crash has yet to reoccur after the power supply replacement on
Feb 24. The stand has crashed once in this period, but was not associated
with a "SCSI error". Please note that a stand crash is a
system/software freeze which is characterized by the microscope stand
motorized controls not responding. If this happens:
- 1. Save your images;
- 2. Shut down the LCS software;
- 3. Power off the microscope stand control box on the floor;
- 4. Wait 15 seconds then turn the stand back on. The stage will
move down a few millimeters, and if the upper limit has been set with your
sample use the course focus up to get back close to focus.
- 5. Restart the LCS software
- 6. Open the image database. Double click in the last image scanned
and use the "Reapply button to restore settings. Note zoom and image
size will have to be manually restored.
- If there are still problems try shutting down the software and waiting
until it has shut down, then power off the scanner (switch #5), wait 15
seconds, then repower the scanner, wait 15 seconds then restart the
software, or contact us.
- Feb 24: Leica SP2 confocal - Main power supply has been replaced.
We hope that this will have eliminated the system crashes that are
associated with a SCSI error report. Note that rare stand crashes
- Feb 3: Leica SP2 confocal - Scanhead SCSI controller was
replaced. Firmware reloaded. The system should work fine as of 6
PM Feb 4. But, since the problem is intermittent, we can not be sure
that the repair effective until the system is run for some time.
Therefore we users to make use of the system. But please let us know
of problems are encountered.
- Feb 1: Leica SP2 confocal - Intermittent problems still persist.
Next scheduled repair is Wednesday, Feb 3. We apologize the continued
- Jan 29: Leica SP2 confocal - Is not working. The SCSI interface replaced
on Friday 29 Jan.
- Jan 14: \\Minsky file server - File shared have been restored.
However errors and omissions may have occurred, so please do not hesitate to
contact Michael or Neal about access issues.
- Jan 11: \\Minsky file server: The computer which is Minsky has
been restored. Shares are slowly being restored. If your data is
not accessible please contact Michael after hours or Neal during regular
hours to have your shares made available.
- Jan 11: \\Minsky file server: Some time between
1:00 AM and 8:00 AM the Windows 2003 server hard disk drive failed. If you needed access
to your data you could have phoned or e-mailed Neal or Michael to
arrange access to the backup server \\Marvin. Since the
Windows system drive is separate to the User Data areas, no user data has
been lost. However the computer which shares this data needed to be
- Nov 3: Leica SP2 Confocal - SCSI interface and micro controller
were replaced. Machine is in shake down mode.
- Oct 27: Leica SP2 Confocal - Again the SCSI VI issue. Repaired
Oct 5 by Dana.
- Oct 20: MHmicroscopy Web server - At 9:30 to10:30 PM Primary webpage
(microscopy.unc.edu a.k.a. mhmicroscopy.med.unc.edu) was taken off line
in order to restore operation from main hardware. All web
pages and calendar pages should be back to normal.
- Oct 19: MHmicroscopy Web server failed after a Windows
security patch was installed at 1:30 PM. Primary webpage was
restarted off a backup server at 2:45 PM. Secondary web pages
(physiology.med.unc.edu, malus.med.unc.edu, confocal.med.unc.edu,
152.19 58.212) remained unavailable until 10:15 PM.
- Oct 14: Volocity License server is temporarily being run
off \\Minsky. Volocity license server status page is temporarily
http://minsky.med.unc.edu:15003 and the connection address for obtaining
Volocity licenses is minsky.med.unc.edu:15002
- Oct 13: Volocity License server is off line due to the failure of
the \\AI computer.
- Oct 13: AI the backup file server is off line due to a power
supply failure and associated failure of both its operating system hard disk
drives. This means that the Volocity server is also off line. No
user data was lost from the RAID drives attached to this server.
However ever daily backups will not be performed while AI is off line.
- Oct 13: Minsky the main file server
was offline from Oct 13 11:45 PM to Oct 14 1:30 AM due to a failed hard
disk drive. No user data was lost!
- Oct 05: Leica SP2 Confocal
Microscope has been serviced .
- Oct 02: Abbe backup computer & raid system - The SCSI disk
controller interface is having issues. No user system is
- Sep 29: Zernike Image Processing Workstation - has been
- Sep 26: Zernike Image Processing Workstation - Is
unavailable due to a burnt out power supply. Impacted - no 64 bit Volocity. Please use
\\Nomarski or \\Snell or your computer instead for Volocity.
- Aug 18: Zeiss 510 confocal - The 543 nm Green laser has
been repaired. All three lasers (blue, green & red) are one
again functioning. Thanks Charlie!
- Aug 14: Zeiss 510
confocal - The green laser (543 nm, Rhodamine, TexRed,
CY3, etc.) has failed. It is flickering and producing
horizontal bands of intensity fluctuations. Repair is scheduled for Aug 18.
- Aug 12: Michael's furlough day part 2.
- Aug 09 - HP1320 b/w laser jet printer is
- Aug 10 - Xerox dye sublimation color printer is
- Aug 08: Printers - Xerox 8550 dye sublimation color
printer and the HP laser jet printer are not shared on Minsky yet
(oops, Michael erred). These printer shares were overlooked
when the Minsky hardware and operating system were upgraded.
- Aug 07: \\Minsky file server - New hardware and operating
system (Windows 2003 server) has been put
into service. All current users data have been migrated to the
new computer. There is no change in the way users should
connect to file shares using \\minsky except for those who use the i.p. number for
access. The i.p. number for accessing \\minsky is now \\18.104.22.168 and also
\\minsky.med.unc.edu. This change is permanent.
(\\22.214.171.124 & \\mh-microscopy.med.unc.edu is no longer in
service.) However there remains the possibility that Michael mis-entered the sharing information for some users/labs/groups.
If users encounter this situation, or other problems, please rest assured that your data
is still stored on the system and is merely inaccessible.
Please feel free to call Michael if you have problems connecting or
can not find shares or files. Please feel free to call
843-3268 almost anytime or leave a message.
- Aug 04: \\Minsky file server - Due to an impending failure the
computer will be replaced. Please note that user data will be preserved.
Yes, the computer event log is indicating that the machine will fail
totally soon. While a replacement computer is
running in parallel access to Minsky may be strange for a few
days. Please call Michael at 843-3268, almost anytime is fine, if you encounter problems.
Aug 04: Zeiss 510 confocal - X-axis scanner
galvo motor has
been replaced. These is no longer a need to scan with 90
degree rotation. Also the galvanometer controller card has
- July 20: Laser Micro-Dissection system has been
taken off the network. Please use an USB drive to transfer
July 16: Michael's furlough day part 1.
Don't ask since I won't tell.
July 16: LeicaSP2 Confocal - UV lens drive
was adjusted. This should fix problems
described on July 3. This problems does crop
up again please let Neal or Michael know so that we
can get the issue addressed.
July 03: LeicaSP2 Confocal - UV lens motor is jamming giving
weak and fuzzy DAPI, Hoechst, etc. images. Work around - try
restarting LCS software. Service request has been submitted.
April 09: Zeiss 510 Meta Confocal Microscope -
Power supply for the Green (543 nm) HeNe laser
was replaced. All three lasers are
April 07: Zeiss 510
Meta Confocal Microscope - The Green (543 nm)
laser has been restored to operation at the expense
of the Red (633 nm) laser. This was done since
many more users rely on the 543 nm laser than the
633 nm laser, and the red and green laser power
supplies are interchangeable.
April 07: Zeiss 510 Meta Confocal Microscope -
The Green (543 nm) HeNe laser has failed.
Looks like this is due to a failed power supply.
Zeiss service has been contacted.
April 02: Nikon upright microscopy (Hooke
workstation) - Power power switch failed.
Switch was replaced April 3. Power connector
to lamp bayonet was also loose. One fault,
with two causes! This was tough to diagnose.
28: Acquisition and Image Processing
Working Stations were turned off Saturday
afternoon in honor of
Earth hour. Files servers will remain running. If you
would like (quick) access to data on Acquisition &
Image Processing computers please telephone
Michael or if you are on site just turn the
computer back on. You will not be forced to
use a candle for microscope illumination.
- Mar 28: Snell - Zeiss 510 full off line software license
has been moved back onto the \\Snell image processing workstation
- Mar 27: Zeiss 510 - The automated
condenser was replaced with a manual unit. Please manually
choose DIC II for Nomarski DIC with the 10x objective and DIC III
for all other objectives capable of performing DIC. Also the
condenser aperture needs to be manually set to minimum for sample
location in order to get maximum depth of focus and contrast, and to
maximally open for DIC and laser scanning transmitted light imaging.
- Mar 26: The Minsky file server small raid's
bad hard disk drive has been replaced. Write access will still be
disabled on \\minsky\users and \\minsky\users1 for now
until the computer is replaced. Please keep on using alternate storage
location(s). e.g. \\minsky\users2
- Mar 26 - Minsky the file server was temporarily unavailable from
5:30 to 6:45 PM while a replacement hard disk drive was being resynchronized
in the small RAID 5.
- Mar 23: Zeiss 510 - The automated
condenser will be replaced with a manual unit. This is being
performed to see if system lockups will be reduced. After this
has been done users will have to manually adjust the condenser
aperture and ensure that the correct Wollaston Prism for DIC has
been selected i.e. II for the 10x objective and III for all other
objectives when doing DIC transmitted light imaging.
- Mar 18: The Minsky file server - small server raid is close
to failure. Do not store data in
\\minsky\users1 This shared directory has been made
read only. Please store new files in an alternate location.
e.g. \\Minsky\users2 or \\Minsky\your_lab-name. Ask
Michael or Neal for assistance and requests for share setup for your
lab or you. The failure of this system is unavoidable since
the hard disk drive in this RAID has been discontinued, and the
replacement drives received were not compatible. If and when
the failure occurs we do not expect data loss, and only minor
disruption while the backup system is switched in. Please note
data older than 24 is backed up on \\Marvin. As part of this
switch over \\Minsky's i.p. number will be changed to 126.96.36.199.
This will only affect those accessing the files server from outside
of the Med. Sch. who access \\Minsky using
- Mar 18
- Minsky small RAID impeding
failure. Write access to \\minsky\users & \\minsky\users1 has been disabled. For the very
small number of users who still use these shares please use another location for
new data. This has been done in order to avoid loss of
data if the RAID array fails.
- Feb 17: Zeiss 510 confocal computer was replaced with a
faster, bigger disk, more RAM computer. Zeiss 510 confocal user
profiles will need to be transferred in order to have filter,
setup, scanning, etc. parameters available on your log in on the new
computer. The new computer is
call \\Hertz and the old computer is call \\Gauss. The old computer
(now called \\Gauss) will remain in service for a few
months, or until it fails, so users can retrieve their Zeiss
profiles and access files. Files on the retired computer can
be accessed using you MHmicroscopy user name and password.
- Jan 21: Spinning Disk Confocal: Nipkow - now has a
transmitted light shutter. C-Imaging users will need to have
its profile updated in order to drive this shutter.
- Feb 06 - Zeiss 510 confocal computer is failing
and will be replaced Feb 17. Also stuck condenser will be
- Feb 01 - File server AI boot hard disk drive has
failed. Daily backup will not be performed until this system
is restored. This is of little consequence to users, since data
is stored on a separate RAID array which is not affected.
- Jan 12 - Hamamatsu OrcaER cameras have been send
out for interface and
firmware upgrades. Affected stations:
Morpheus (time lapse &
Fura-2) - Nipkow (Spinning disk). Expected return: early Feb.
- Jan 06 - \\Marvin has been replace with new
- Jan 03 - Backup server \\Marvin has failed
- Dec 28: Image Processing Workstation called \\Brewster has
failed. \\Nomarski has replaced Brewster.
- Oct 14: Minsky, the file server - A drive controller and its
associated hard disk drive catastrophically failed after 4:00 AM on October
disk partition serving Lamantia, Faber & Blancafort labs. and
Dentistry, Dermatology & Psychiatry Departments shares were totally
lost. This unit was replaced on Oct 13 and files restored from primary
backups on Oct 14. As far is I can determine no data file was lost. Please
Michael if you encounter any outstanding issues.
- Aug 28: Zeiss 510 confocal - 633 nm (red)
has been repaired.
- Aug 28: Zeiss 510 confocal - The red (633 nm) laser is not firing
up. Zeiss service has been contacted. Imaging with the 488 nm
(blue Argon/1) & 532 nm (Green HeNe) lasers can still be performed.
- Aug 22: Microscope stand on the Leica SP2
Confocal has been reset. Focus and limit setting is working
- July 24: Leica SP2 Confocal is down. Systems is running again.
Scanhead microcontroller was replaced.
- July 18: Leica SP2
Confocal is down. Part is on order and expected the week
starting July 21.
- July 17: Leica SP2 Confocal is down. Error reported is "Timeout in device: SCSI Controller V1".
The machine can not be used for scanning. Leica service has been
notified, and we are waiting for a scheduled repair time and date.
- July 08: Leica SP2 Confocal - Scan head voltage adjusted to 5.6V
- June 26: Zeiss 510 Confocal - X-Galvo replaced.
- May 29: Leica SP2 Confocal has been repaired
- Power supply to scan head was readjusted.
- May 28: Leica SP2
Confocal is still down. Parts expected to be here
today did not arrive. They should be here tomorrow. If
you do not need UV, the Zeiss 510 may be an option - contact
- May 21: Power failure: Floor buffing caused a circuit breaker to
be tripped from 10:50 PM to Thursday 12:25 AM. Unavailable servers \\Minsky
and secondary web pages, \\Marvin. Some daily backups were interrupted.
Dear Bill Marzluff, please we need some server room space!
- May 19: Leica SP2 Confocal is down. Local
service technician is out of town until Friday but we will try to get Leica
to send someone sooner.
- "Timeout in device: SCSI Controller V1"
- Update May 20:
- Talked to Miriam in Tech Support at the 866 number walked me through
some restart procedures and we confirmed that there's nothing we can do
here. A job order has been opened with Leica.
- Dana, our local service tech, will be here Tuesday to work on the
system. He is currently out of town on service calls.
- UPDATE MAY 28
- Because of the stalled PO for the service contract, Leica service did
not send the requested parts to Dana. Dana has requested the parts be
sent today based on our previous verbal commitment for the service contract.
He will plan to come tomorrow once the parts arrive.
- Apr 14: SERVER \\MINSKY
IS RESTORED Logons and access to shared directories and
files are restored. Please contact Michael if you notice any
- Apr 14: SERVER \\MINSKY IS DOWN
Went offline this morning. We are in the process of getting it
back on-line and hope to have that complete before the end of the day.
File access and logon to MHMF computers are affected. See a
facility director if you need to log on to a computer.
- April 07: Zeiss 510 confocal - AOTF control unit cooling fan
- Mar 26: Zeiss 510 confocal - Reported to Zeiss central funny
sound from laser box at startup (probably a fan) and increasing
amperage at consistent power output %. Work order # 41037839.
- Feb 28: Leica SP2
AOBS Confocal -
working. PMT preamp and power supply board were
- Feb 19: Image Processing Workstation \\Hooke - Replaced dead 21" monitor with used 19" CRT.
- Feb 15 2008: Leica SP2 AOBS Confocal - Service Tech visit.
- Replacement of PMT#3 did not solve the problem.
- Connected PMT#3 to new port on controller board and had it working
briefly but failed again before he left. The system is still
functioning with only PMTs 1 and 2. Will order new controller board.
- Some correction of UV illumination evenness.
- Jan 04 2008: Leica SP2 AOBS Confocal -
PMT #3 is not working properly. To get any signal, offset
must be turned to 40 or greater. Please see a
facility director for
assistance in adjusting your acquisition parameters. Service has
been contacted. We were notified that there is a long line ahead of us
so it may be several weeks before this is fixed. In the meantime:
- PMT #2 should be used instead of PMT#3 for all imaging parameters.
- Many of the Beam Path Settings for acquisition will need to be
- For imaging more than 2 channels, sequential scanning with frame
switching will need to be used.
- UPDATE Feb 11: Our Leica service technician is scheduled to come this
Friday to address the PMT3 problem and the uneven UV illumination with the
- Dec 04: Zeiss 510 Confocal - The D: drive
was failing, and the C: & D: partition hard drive has been replaced.
Due to file corruption on the original D: drive most files were not restored
to it. However all C: drive files and the operation system were
restored. Failure was progressive and it is not certain when this
problem started. Data stored on this drive was last copied on Dec 03 and can
be made available upon request to Michael.
- Oct 14: Raid File Server \\MARVIN - The Windows 2003 operating
system for the computer called Marvin has been rebuilt from scratch.
Raid Server Data is present and intact. If you use this server for
file storage you will notice that there are currently no user network shares
available (shared directories, e.g. \\marvin ). If you do use these
shares on Marvin please let Michael know the share name you used, so that
sharing can be re-established, and your data made available.
- Oct 10: Snell image process workstation - Software installed:
C-Imaging, Zeiss LSM 510, MetaMorph 7.1, Adobe Creative Suite 3, WinZip 11
|Zeiss AIM 510 (full off line)
|Leica SP2 (full off line)
|QuickTime Pro 7.2
|Adobe Creative Suite 3
|Roxio DVD/CD writing
|Symantec Anti Virus 10.1.6.6010
- Oct 9: Raid File server \\MARVIN - Is not
booting up after Microsoft's update Tuesday. Data should
be secure on the raid array, but the computer which coordinates access to
the array is in a constant loop of rebooting. Tests are being carried
out to determine what is bad component in the system. Components found
to be good so far are: CPU & memory, RAID array drive,
- Oct 4: Snell image process workstation - Is available
again, but software still needs to be reinstalled. Data on
hard drives is preserved. Files that were left on the Windows
desktop can be found in N:\Documents and Settings\ in a
directory with your username. i.e. N:\Documents and Settings\%username%\
or N:\Documents and Settings\%userprofile%\desktop
- Oct 3: Volocity - Is now on the
Windows x64 bit workstation call \\Zernike which is currently located on
the large table in 6129 Thurston Bowles. It will run faster in this
machine. The D:\ drive is shared as users1. i.e.
- Sept 13: Snell image process workstation - Is not available
due to motherboard & Video card failures. Data on the hard drives will
be preserved, and the computer may be back on the network on Monday.
Software will be reinstalled subsequently.
- August 17: Leica inverted
- Fine focus has been repaired. A deteriorating FITC cube has
- July 26: Nikon Microphot - cleaning, lubrication. All
transmitted light filters now work.
- July 26: Leica 4D - Is back in operation
- July 25: Leica 4D Confocal - Optics clean, lubrication, adjust
condenser arm microswitch, clean stage, align visible laser, mount laser
back on unit, laser alignment
- July 20: Leica SP2 AOBS confocal - Detector board replaced.
PMT2 channel is much cleaner. UV realigned in Z axis.
- July 18: Leica inverted - An axis for the focus drive is cracked.
Course focus is unaffected but the fine focus may have slight delay.
Replacement parts are on order.
- July 18: Leica inverted - A new transmitted light power supply has been
installed. Transmitted light control is back to normal.
- July 5: Leica SP2 confocal - Alignment of all channels was
performed in XY and in Z.
- June 29: Leica SP2 Confocal -
Files older than 6 months have been
purged from the D: & E: drives
- June 26: Leica inverted -
A temporary transmitted light power supply has been setup. It
has two power levels, off and on - please use the power off /on switch.
The rheostat on the temporary power supply is faulty and will be replaced.
For now please just use it a maximum power only.
- June 25: Leica inverted - Transmission
lamp power supply has burnt out. A temporary power unit will
be setup on Tuesday. Fluorescence viewing and imaging is unaffected.
- June 25: Leica Inverted Microscope:
File older than 6 month on \\Fresnel\users2 (the computers d: drive)
area have been purged.
This action is in accordance with
facility policy, and was carried out due to no free space
on \\\FRESNEL\USERS2. Removed files may be
present on alternate storage. Please contact Michael for
- June 25: "Roach" raid array: Left redundant power supply is being
flagged as failing. PAM is not responding. However file shares
on \\marvin are still accessible. Note to self: Make sure the mobile
phone is well charged and call Promise.
- June 22: Minsky server - Files older than
2 yrs have been purge from \\MINSKY\USERS2.
This action is in accordance with
facility policy, and was carried out due to limited available free space
on \\\MINSKY\USERS2. Removed files may be
present on alternate storage. Please contact Michael for information.
- June 19: Nikon TE2000 MultiMode - The DG-4 fluorescence
illumination is uneven and not as bright as it should be. Technical
assistance has been contacted.
- June 15: Image Processing Snell - The power supply was
replaced and Snell and is working well so far. Please alert
a facility director if you have any problems.
- June 11: The image processing workstation
called Snell is crashing and failing to boot. Cause is
being investigated. A marginal power supply is suspected. i.e. too many
devices inside this computer (3x hard disk drives, high performance video
card, parallel card, DVD writer) just exceeding the maximum power which can
be delivered from the existing power unit.
- May 14 Leica SP2 Confocal:
The UV laser excitation path has been realigned and should be
working. Please contact a facility
director with any problems.
- May 03 Leica SP2 Confocal: The system is functioning again!
- May 02 Image Processing Workstation \\Snell: The computer
has been locking up occasionally. This seems to e hardware
related. DVD/CD burner has bee replaced since it has been
unreliable. These problems may be due to an inadequate power
supply. This system runs with 3 hard disk drives and a
DVD burner off a 400 W power supply. A larger power supply
will be tested in this system.
- May 01 - Leica SP2 AOBS Confocal: Replacement boards
x-y-z galvo and reassigning the Mic Stand control to COM2 allowed
the system to run for ~30 minutes, until a board failure brought
down the system.
- Apr 27 - Leica SP2 AOBS Confocal: Replacement of additional
galvo components (control board) and 4 hours of troubleshooting did
not solve the problem. More parts will be ordered and another
round of troubleshooting will begin on Wednesday, May 2.
- Apr 26 - Leica SP2 AOBS Confocal: Replacement of the
x-galvo components that arrived today did not solve the problem.
Now we must wait for parts sent from Germany.
- Apr 25 - Leica SP2 AOBS Confocal: The
system is non-functional due to the failure of the x-galvo
mirror for laser scanning. Replacement parts are on
order and should be installed Thurs. 4/26/07.
- Apr 12 - Zeiss 510: HRZ/Galvo stage is working properly.
Transmitted light switching mirror was replaced which should prevent
further hardware communication errors.
- Apr 12 - Zeiss 510: HRZ/Galvo stage is not
responding. X-Z scanning can still be performed slowly with "z-sectioning"
i.e. stand fine focus control. However x-z scanning with the
high speed hyperfine mode is not possible. Note that standard
z-series scanning is still possible.
- Apr 3 - Zeiss 510 confocal: The hardware
communication error for the transmitted light mirror continues to
appear intermittently. If you encounter this and are unable to
view or image DIC, close the software (lasers stay on), power cycle
the microscope stand and restart the software (Argon laser power
will need to be re-set).
- Mar 30 - Zeiss 510 Meta confocal: The microscope
stand power supply unit and the transmitted light PMT cable were
- Mar 09 - Zeiss 510 Meta confocal: The condenser unit was
- Mar 8 - Leica SP2 AOBS confocal: The
microscope fluorescence module was adjusted. Please alert
a facility director of any further stand crashes.
- Mar 07 - LeicaSP2 Confocal: Stand service, especially
fluorescence cube turret and shutter electronics. Replace
lower AOBS AOTF controller fan.
- Feb 16 - Zeiss 510 confocal: The transmitted light
mirror is sticking intermittently. Service has been contacted
and parts on on order..
- Feb 15 - LeicaSP2 AOBS Confocal: The microscope stand is
continuing to crash periodically. Service has been contacted.
- Feb 15 - Web server: The computer
serving the pages listed below will be upgraded soon. This
change has no impact except for users accessing the FTP site and
those administrating on the calendars. A note will be placed
here when the site is upgraded.
- Feb 12 - LeicaSP2 AOBS Confocal: Original microscope stand
controller put back into service, and firmware pack 4.3.1 loaded.
- Jan 17 - Image Processing Workstation Snell: Defective
- Jan 17 - Image Processing Workstation Snell: Drag-Direct-to-DVD/CD removed since it was slowing down the system
significantly. Logging out should now be swifter.
- Jan 08 - LeicaSP2 AOBS: Several bad crashes due to
unpredictable stage behavior. Reset parfocal settings for all
objectives and seems to be working better. Cause of
parfocality disappearance unknown.
- Jan 02 - Nikon Inverted TE2000 -
The DVC 1412 camera is working again. A p ower supply unit was
faulty. A replacement power supply has been put into service.
The DVC 1312 camera is still connected to Morpheus and may be used.
- Dec 24 - Nikon Inverted TE2000 - 1412
camera is not functioning. 1312 is now connected to the
computer called Morpheus. Problem will be appraised in early
- Nov 17 - Minsky:
Another hard drive has failed on the file server Minsky (a different
unit to the Oct 31 failure listed below). It was replaced and restored
with data prior to Nov 11, and was back in service on
Saturday morning. Affected
are listed. Note: since files were restored from cumulative
backups, file which have been deleted may be back in user
- Oct 31 - Minsky: A
failed hard disk drive has been replaced. Affected files have been
loaded from backups made up to Oct 14, 2006. Affected users are
listed in this link:
stored on the failed drive from Oct 14 to Oct 28 may still be
present on the original computer workstation. If the files
have been deleted from the workstation they may still be available
on separate workstation backups of Oct 17 & 24. Please contact
a facility director if you wish to examine these off line backups.
- Oct 28 - Minsky: A drive has failed affecting shares and data of
- Oct 25 - Leica SP2 AOBS Confocal:
The UV light path was aligned for better image quality.
- Oct 24 - Color Printer: The Xerox Phaser 860 color
printer is experiencing error 15,004.45:8029 and is not available for
- Oct 23 - Zeiss 510: Galvo Z-stage recalibrated.
Please let facility personal know if you notice instability or non
linearity in the galvo stage z-positioning system.
- Oct 21 Sat. 7am-10am:
Scheduled Power Outage for Thurston Bowles
- Ensure all lasers and fluorescence light
sources are powered down prior to the power outage.
- MHMF equipment will be unavailable during this
- Servers will be shutdown at 6:00 PM on Friday
and restarted Monday morning (please let a
facility director know if
you need access before Monday)
- Oct 20 - Leica Inverted Fluorescence Microscope: Imaging
of far-red emitting fluorophores (ie, Cy5, AlexaFluor 633,
AlexaFluor647) is now available using a new filtercube. See
a facility director for instructions on how to implement this new
- Oct 12 - Minsky file server: Many shared
directories were unavailable due to two disk packs disconnecting from
the computer. Data on the drives was safe, but unavailable on
the network. The network shares were
restored at 2:30 PM Oct 12
- Oct 12 - This web server: Down - Due to a waste
paper bin placed against the key board this server did not
reboot after a Windows security patch installation. Down 11 PM Oct
11 to 9:10 AM Oct 12.
- Sept 22 - Marvin file server: is available again.
- Sept 25 - Zeiss 510: confocal is back in action. A scope
interface board, the condenser and transmitted light changer were replaced.
- Sept 22 - Marvin file server: is not
secondary file server, Marvin is failing to boot up. The data on it is
safe and will be available again once the operating system for Marvin is
repaired or reinstalled. Please note that most users access the main
file server called Minsky. Only a minority of users will be affected
by the unavailability of Marvin. If you urgently need data on Marvin
- Sept 20 - Leica SP2 confocal:
Frequent microscope stand crashes. All motorized controls fail
to respond to button presses, focus is frozen on the stand, and LCS software
will not run microscope (e.g. switch from scan mode to visual mode).
- To fix problem:
- 1. Save images
- 2. Shutdown LCS software -
- 3. Allow LCS software to close then
Power down microscope stand controller
- 5. Re-power microscope stand
- 6. Restart LCS software
- 7. Please record "Stand Crash" in log book
- Sept 20 - Zeiss 510 confocal: Scanning is
occasionally locking up.
Service has been scheduled.
- Sept 5 - Leica TCS-4D: The Kr/Ar
laser would not ignite this morning. We will work on this as soon as
possible. In the meantime, the Zeiss 510 has significant availability
- Aug 24 - Servers Minsky & Marvin: An
uninterruptible power supply (UPS) battery failure forced the
partial powering down of the file servers Minsky & Marvin.
Some users' shares and printers on Minsky and Marvin were
unavailable for ~12 minutes from about 10:15 AM.
- Aug 16 - Leica-LMD: Preventative maintenance service;
included realignment of laser for better power.
- Aug 16 - Inverted Leica fluorescence: Service was performed
on the stage to address the problem with lateral drift.
- Aug 15 - Leica SP2 confocal: A borrowed 40X 1.25NA oil
objective has been put on the system while our objective in place of ours.
An improvement in the blue channel should be noticeable. Our objective
has been sent out for repair.
- Mel Gibson was born in New York: He only lived in Australia, and
studied at the University of New South Wales, Australia.
- Aug 1 - Zeiss 510 confocal: The
F: drive storage space has been substantially increased to 110 GBytes.
- 1. Files on the F: drive older than May 1 are no longer present. (In
effect files older than May 1 have been erased, but don't panic and see
point 2. below.) This is because a new F:
disk drive has been installed in the system, and only files newer than three
months old have been copied to the new drive. Therefore problems may occur
if you are opening image databases from the F: drive (or copied from
the F: drive after Aug 1) that have images
saved prior to May 1, 2006.
- 2. However for a limited time (until Aug 1, 2008) files and image
databases that were present at Aug 1, 2006, can be made available by
- Yes, this is all rather complex. However we do not want to keep
"expired" data, since both the D: & F: drives do
fill up. In fact we spend a significant amount of time dealing with
expired data, and hope users appreciate that it is difficult for us to
determine which are
unneeded files. It should be noted that most facilities have draconian
rules about old data. We know that a large fraction of the files on
the image acquisition systems are well backed up by users. However we
have very little idea which files are actually no longer required. In
some ways we create this problem by advising users not the erase data files
until they know the files are securely and adequately transferred to their
computers/servers/CD/DVD/etc. But transfer errors do occur
bottom line please see the facility file storage policy.
- Those who store image databases/files on the D: drive may
consider using the open space on the F: drive instead.
- Aug 1 - Zeiss 510 confocal: Scanning is
back to normal after the service visit. Scanning galvos were
recalibrated. Swing arm safety switch repaired. Annual PM
performed. F: drive upgraded to 110 Gbytes from 39 GBytes.
- July 31 - Zeiss 510 confocal: Scanning is
intermittently failing. Service will be carried on Aug 1 in order to
investigate the cause. Check back here for updates.
- July 31 - File server \\MINSKY:
\\MINSKY went off-line starting at 1:24pm and was back
on line at ~2:00pm. Access to files on the server and MHMF computer
network drives were affected.
- July 10 - Zeiss FRAP demonstration: Elise
Shumsky of Zeiss will demonstrate how FRAP can be performed on the
510 Meta confocal microscope. Let a facility
director know if you plan
- July 4 -
Michael Chua: I
am receiving spam e-mails faster than I can erase them. If you
need to contact me please use the phone, paper or other 19th century
and before devices
- June 7 - Zeiss 510 Meta confocal: The mercury lamp power
supply was replaced with a new unit. Start up and shut down procedures
are the same as before.
The HRZ stage continues to misbehave. The Zeiss service technician has
ordered a new stage.
- June 6 - Analysis Workstation \\Young: Replaced failing video
card with MSI nVidia FX5200
- June 2 - Leica Inverted Fluorescence microscope: The stage
has been tightened to prevent drifting. Movement is still not
perfectly smooth. Please alert a facility
director if you continue to
have drifting problems.
- May 31 - Image acquisition station \\Hooke: Replaced video card
because of a failing fan on the previous card.
- May 31 - Leica Inverted Fluorescence microscope: The stage
is drifting when it should be stationary. The service engineer will be
here Friday, June 2 to address this issue.
- May 30 - Analysis Workstation \\Young: A cooling fan on the
video card is making a rattling/humming noise. The computer still
- May 23 - Zeiss 510 Meta confocal: The service engineer
could not replicate the z-galvo stage problem, so it is still unsolved.
Please record any troubles in the log book and alert a
- May 12 - Zeiss 510 Meta confocal: the
z-axis galvo (HRZ) is not reliably returning to the beginning and end set
positions. Occasionally a buzz can be heard emanating from the HRZ
stage transducer box. Zeiss service has been informed. Engineer is expected
about May 19. Temporary fix is to try powering down the system and
- April 12 - Leica SP2 AOBS confocal: The microscope stand
controller was replaced with a new unit. Please alert a
if you experience any troubles with the system.
- April 6 - Leica SP2 AOBS confocal: A
hardware communication failure may occur in the middle of your scanning
session. The service technician has been contacted and is working on a
solution. If you experience a hardware communication failure:
- Save your work
- Exit the Leica LCS software
- Turn off the Microscope Stand Controller (#2) (ONLY the microscope
stand, no other components need to be adjusted)
- Wait 15 seconds, then turn the Microscope Stand Controller on again
- After the microscope completes startup, restart the software
- April 6 - Zeiss 510 META confocal: The transmitted light
mirror was replaced.
- Mar 31 - Zeiss 510 Meta confocal: A
"Hardware is busy" message may appear when scanning is initiated.
Transmitted light mirror motor is jamming and will be replaced in early
- Mar 28 - Minsky file server: At 3:07 PM a transient
drive outage occurred due to a failed cooling fan. The drive was
back on line at 3:40 PM. Some shared directories were affected.
- Mar 20 - LeicaTCS-4D confocal: is back in
action with a functioning laser! In order to get the system
operational the PC computer was replaced twice! The current
computer has the same disk drive as \\Jansen (the computer installed June
2005), and hence the network identity is unchanged and user image files are
still there. Laser light power at the stage is good. Polarization of the
laser has not been fine tuned yet, so DIC/Nomarski may not be optimum for
now, however fluorescence is not compromised.
- Mar 20 - Workstation computer \\Brewster has replaced
\\Monarch. Files from \\Monarch were backed up to a protected area
of \\Marvin prior to the exchange. Please see a
facility director if
access. Monarch has been decommissioned, and software reinstalled on
- Mar 20 - Minsky file server: Some shared
drives on the server \\Minsky were unavailable
for a short time today due to a failed power supply and seized cooling fan
on one of the firewire drives. Access to all drives has been restored.
- Mar 17 - LeicaTCS-4D: Argon/Krypton laser has been received and
is being installed. The computer for this system is not starting.
- Feb 27 -- Spinning Disk Confocal: Installation and testing
of new filters and correction lens has been completed.
- Feb 27 -- Inverted Nikon MultiMode: Phase Contrast ring for
the 4X objective has been added to the condenser turret.
- Feb 22 -- Zeiss 510 confocal: Mercury Arc lamp is not
igniting on the first attempt. Please make sure the round LAMP
indication light (NOT the power switch light) glows green before turning on
the Remote Control power. If the lamp does not ignite, turn off the
power switch, wait a few seconds and try again.
- Feb 20 -- Zeiss 510 confocal: Loss of
communication between the software and microscope stand is occurring
intermittently but with increasing frequency. To restore
communication: 1. Exit the software, 2. Reboot stand, 3. Restart the
- Feb 18 -- Zeiss 510 confocal: A failing AOTF fain in the laser
module is creating a groaning sound from under the lasers when the software is started.
Service has been contacted.
- Feb 13 - Zeiss 510 confocal: A cooling fan has failed in the
power supply of the computer. It is possible that the computer will
shut down without warning if the room temperature gets too warm.
- Jan 27 - Laser Micro Dissection: Continuous erratic UV laser
firing was due to trigger being incorrectly set to internal. Switching
back to external trigger returned laser firing back to normal.
- Jan 26 - LeicaSP2 confocal: UV laser seems to working better once
warmed up. Poor performance may be due to a bad mirror alignment.
An automatic realignment will be attempted.
- Jan 24 - LeicaSP2 confocal - 40x oil x-z image quality is not as
good as expected for blue emissions light (such as DAPI). 40x Oil
objective cannot be used for z-axis co-localization with blue emitting fluorophores.
Red and green co-localization is working great. (UV co-localization is
good with the 63X 1.4 AN oil & 20X objectives.)
- Jan 24 - Leica SP2 Confocal - UV laser
behaved erratically. Please alert a facility
director of any erratic behavior you encounter.
The UV laser
failed on Jan 24, and was only operating at very low output power. In
such circumstances, it can still be used with appropriate power adjustments.
1. Turn power knob on the control box
1/4 to 1/2 turn; be sure the "Current Limit" light is not illuminated.
2. Adjust the 364nm AOTF power as
high as necessary (up to 100%) for a good image at Gain setting of 500-650
We are working to diagnose the problem. Please stay tuned for updates.
Update: 5:30pm. After running at minimum power for
several hours, the laser returned to normal behavior. The power
successfully increased to 2 full turns of the power level knob 2:45pm and
continued to run normally for several hours until it was powered off.
- Jan 06 - Fresnel - The hard drive has
been replaced and is available for imaging. IMPORTANT: All
users must run the IMAGING_SETUP.reg file before they can use SimplePCI.
Before starting SimplePCI, double-click on the desktop icon for
IMAGING_SETUP.reg. Answer "Yes" or "OK" to the 2 dialog boxes that
appear. If there are no cameras available in SimplePCI, exit the
software and run the IMAGING_SETUP.reg file from the desktop.
- Jan 05 - Fresnel -
The hard drive on the computer Fresnel has been replaced and needs to be
reconfigured. The system is still available for viewing samples but
imaging is not available.
- Jan 01 - Minsky shares for some labs. and
users: will not be available on the network until a hard disk drive on
the Minsky file server is restarted on Jan 03.
Affected shares are listed below. This problem arises due to a not
well documented problem with large data transfers via Firewire interface
controllers. Some users will have their share or their lab. share on
Minsky mapped to to drive
letter (typically m:) on various work stations. If this mapping
connects to a share listed below the drive letter will not be available.
||Shared Path (Hidden to users)
Notes: For administrative purposes this drive is called users3,
although users do not have access to this drive though the \\minsky\users3
- Dec 13 - Xerox color printer:
Maintenance kit has been received and installed. Color printer
is back in action.
- Dec 12 - Xerox color printer:
Printing is not possible until a new roller is received. This
roller is called a "maintenance kit" by Xerox and is in fact a
roller impregnated with silicone oil, which aids in the release of
the paper from the print drum. This kit has a chip
which inhibits the printer from printing from 366 days after it is
1st used, so printing can not be forced. A replacement was
ordered, and we are still waiting on delivery.
- Dec 06 - Leica 4D confocal: Triple line
laser is with Melles Griot for repair
- Dec 01 - Leica 4D confocal: Triple line
laser has been shipped out for repair.
- Nov 04 - LeicaSP2 confocal: UV fiber
alignment has improved power the level to the stage, and lens alignment has
returned x-z imaging quality to expected levels for the 63x oil, 63x water,
20x, 10x objectives. 40x oil x-z image quality is not as good as
expected. User will note that required 364 nm AOTF levels are
back to normal.
- Oct 25 - LeicaSP2 confocal: UV sensitivity
and z-axis alignment has been improved on the 63x, 40x, 20x & 10x objectives.
Note: run the 364 nm AOTF higher rather than increasing the UV laser tube
current. UV z-axis imaging still exhibits some dispersion, however x-y
alignment is still good. This dispersion issue will be addressed at a
date yet to be determined.
- Oct 25 - Marvin - secondary file server:
Due to what can only be described as excessive weirdness OIS suggested
changing the IP number for \\Marvin & marvin.med.unc.edu to 188.8.131.52.
It may take a day or two for this change to propagate to users desktop
computers. This means access to \\marvin or
\\marvin.med.unc.edu may not succeed during this period. As a
work around during this time, or anytime in fact, try connecting by placing
\\184.108.40.206 in the address bar of a "MyComputer" window instead.
- Oct 19 - Minsky - domain controller: A disk drive on Minsky's
primary disk array crashed. From 2 to 4 PM Minsky was unavailable
while the drive was being rebuilt. Network and local logins along with
access to shared directories were not available during this period.
- Oct 10 - LeicaSP2 confocal: A new UV fiber optic and pinhole were
installed, however UV signal remains week and z-axis alignment is only good
with the 63X 1.4 NA objective. Repair is still pending. This
problem remains seemingly obstinate. A new fiber and pin hole will be
installed on Oct 25.
- Oct 05 - Zeiss 510 confocal: Green 543 nm laser and power supply
have been replaced
- Oct 04 - Zeiss 510 Confocal: The Green
543 nm laser has burnt out. Replacement is scheduled for the morning of Oct
05. It is expected that the system will be fully functional again
before noon of Oct 05.
Some users are having trouble connecting to
\\Minsky. If you encounter this situation use its IP number
instead. i.e. \\220.127.116.11
- Sep 1 - Minsky: The drive on Minsky which
hosts the shared directories listed below continues to intermittently not be
recognized by the operating system after intensive reading or writing to the
drive. These directories, as they existed on Sep. 1, have been copied
to an alternate drive system on another computer called \\Marvin
Access to these directories will be setup as quickly as reasonably
possible in the next day. If you required immediate access to files on
these shared directories and \\Minsky\yourname or \\Minsky\lab-name
seems not to be available please contract Michael (843-3268)
- Aug 31 - Minsky: The above lab. and user shared directories
on Minsky are currently offline.
It is expected that power cycling of this drive on Sept. 1 should restore
access to the files.
- Aug 30 - Minsky: Old Minsky has been
restored. The domain is once again called
MHMICROSCOPY (not mh-microscopy). Files shares are back to
the original state except for data stored under \\minsky\users1 when it was
part of new Minsky (mh-microscopy). These files will be transferred
back to \\Minsky\users1 on Aug 30.
- Aug 26 - Minsky:
A new Minsky computer has been created.
Although the computer name is still \\Minsky,
its domain name is slightly different and is mh-microscopy (note the
minus between mh & microscopy)
- Aug 26 - Minsky (domain login controller and server file
shares) - A RAID controller card for Minsky has been ordered.
Delivery is expected on August 31. Updates of the status of the RAID repairs will be posted here.
Also see just below.
- Aug 25 - Minsky the main server & domain controller computer has
a bad RAID disk controller. The immediate consequences are:
- Logging on to any MHMICROSCOPY computer may be impossible with
your user account. Please see a facility
director for assistance.
- File shares \\Minsky\users1 & \\Minsky\users2 are not
- Mac file access to user1 is not available
- Personal and lab. files shares are not available.
- Printing to the color Xerox and b/w HP printers is
- A replacement controller is being ordered from Dell. The worst
case scenario is that the RAID controller can not be replaced, in which case
the files on the RAID drives can not be accessed. This will affect
\\Minsky\users1, and all user accounts will have to be issued new passwords.
Not affected are \\Minsky\users2 & individual and lab. shares of the
form: \\Minsky\yourname and \\Minsky\labname
- Aug 23 - Laser Micro Dissection: The UV laser has been
replaced and is functioning normally.
- Aug 9 - LeicaSP2 confocal: The red 633
HeNe laser has been successfully replaced. The UV laser alignment was
also adjusted. Please a facility director
if you have any questions or troubles.
- Aug 2 - LeicaSP2 confocal: The red 633
HeNe laser has burnt out. Excitation of CY5, DyLight 647, etc. can not
be performed until the laser tube/power supply are replaced.
Replacement will be done on Monday, Aug 8, contingent upon the replacement
units being received. Check back here for any updated information.
- July 27 - Spinning Disk Confocal: The spinning disk
confocal will be sent off September 2 for upgrades to improve field
intensity uniformity and synchronization of the camera with the spinning
disks.. Estimated return date is September 13th if no additional
problems are encountered.
- July 9 - Laser Micro Dissection: UV laser
is flickering between cuts. New laser is back ordered.
- July 5 - \\Minsky\users3: Laboratory and individual shares
on Minsky have been restored.
- June 30 - \\Minsky\Users3: The
drive containing the individual lab folders previously accessible on \\Minsky\Users3 has been recovered.
It is now Drive F: on the computer Nipkow, which
is the computer on the spinning disk confocal. The drive is currently
not shared on the network. Users with data on this drive
should be able to log into the workstation Nipkow and access their files for
transfer to CD or flash drive. We will work to restore network access
after the July 4th holiday if not sooner.
- June 29 - \\Minsky\Users3: The
individual and lab. folders kept on
\\Minsky\Users3 are currently not accessible. The drive enclosure
failed. We will do our best to retrieve the files as soon as
possible. Please contact a facility director if you have
- June 28 - Computer \\Hooke: Old
files from the computer \\Hooke (Nikon Upright microscope) can be accessed
at network location \\Nipkow\Hooke_Users2. Please contact a
facility director if you have any questions.
- June 22 - Nikon Upright/Computer \\Hooke:
Computer \\Hooke will not power up.
Imaging is now possible using
computer \\Jansen. We are still working to
access to files on Hooke's hard drive.
NOTE TO SIMPLE-PCI USERS on this system: All settings have been
reset to null and will have to be re-established after being
retrieved from Hooke's hard drive. Please see a
facility director for help with this.
- June 10 - Zeiss 510 confocal: The fluorescence filtercube
seems to working properly again. Please report any further problems to
a facility director.
- June 7 - Zeiss 510 confocal: The
fluorescence filtercube turret on the microscope stand is not properly
switching between Visible and LSM modes and not selecting the correct
filtercube position. If needed, move the carrousel by hand until it
clicks into the correct position. Before scanning, ensure that
Position 1 is correctly selected (the carrousel needs to be in
this position). If the
carrousel is not correctly seated in Position 1 for scanning, there will be
a 5s pause before each scan and you will get an
Zeiss service has been contacted. Please find a
facility director if you
have any questions.
- June 3 - LeicaSP2 confocal: Measures have been taken to
reduce system software crashes. To help us asses the effectiveness of
these steps please note any crashes you experience in the log book, and
record what you were doing and which keys/buttons were pressed immediately
prior to the crash/system freeze. Thank you.
- June 3 - LeicaSP2 confocal: UV alignment has been checked.
DAPI, Alexa 350, Hoechst, etc. sensitivity should be good, and produce low
noise images. Suggested initial AOTF settings for the laser lines are:
UV 361nm=10% (laser power 2 turns) 488 nm=10% (laser power at 9 o'clock)
- June 3 - LeicaSP2 confocal: The E:
drive has been replaced. Data may be stored there. The
network shared directory \\LAPLACE\USERS2
is also available again. Data that was on the drive will not be
restored. Users will need to create a suitably named folder for
themselves. Additionally, with a suitable donation of payola, users
may request their data from the crashed drive to be restored.
(Astute users may notice that the replacement drive is half the size (35 GB)
of the crashed drive. A 72 GByte drive will replace this smaller
drive. Any user data present will be transferred to the larger drive.)
- May 21 - Leica 4D confocal: The new laser has been installed. The machine is back in operation (for use by CF
Center users). Please note: 1. start the laser with the switch on standby and
2. shut down by switching
to standby. The cooling fan should run
automatically for 2-3 minutes then the power (rocker) switch should be
switched off. The OS9/VME computer should be turned off at the end of
the day by the last users.
- May 16 - LeicaSP2 confocal: The E:
drive has failed. As a consequence the \\laplace\users2
shared directory, which is a mapping of the E: drive also unavailable.
See a facility director if you wish to attempt to recover data from this drive.
- May 13 - LeicaSP2 confocal: The E: drive on the Leica SP2 AOBS
confocal microscope is currently failing. It is expected that this
drive will fail totally.
- May 12 - Image Processing Workstation \\Snell
(6129 Thurston Bowles): is unavailable and disconnected from the network. Due to a
serious security breach the Image Processing Workstation called \\Snell is
not connected to the network currently. Consequently the following
shared directories are unavailable:
- Software unique to this computer is Volocity 3D viewing &
rendering and Zeiss 510 offline confocal analysis. Hardware on this
computer is the DVD+-R/RW recorder. Please note that
DVD-+R/RW capabilities are available on the Image Processing Workstation
- It is expected that the program drive on this computer will have to be
scrubbed (low level formatted). Data drives (D: & E:) will
remain intact when the operating system is reinstalled.
- Due to the security breach it is possible that passwords entered on this
system from March 10 to May 12 may have been compromised.
- If you have questions please contact Michael
- April 14 - PE Spinning Disk: The
Perkin Elmer spinning
disk confocal is up and running again for live cell imaging of GFP-,
mRFP- and Cy5-like fluorophores. The system is now controlled with the
same C-Imaging software package used with the widefield system in the
Facility. Contact a facility director for more information or to set up
- April 13 - Zeiss 510 Meta: Laser
replacement has been rescheduled for the evening of Wednesday April 13.
Replacement was completed at 9:40 PM. Thorough X/Y galvanometer calibration was
performed. Bi-directional scanning should produce a better
- April 11 - Zeiss 510 Meta:
Bidirectional scanning has been recalibrated electronically. This
should improve the coherence of bidirectional scanning. An optical
recalibration may also be necessary, and needs to be performed by Zeiss.
Note that very few users select bidirectional scanning.
- April 08 - Zeiss 510 Meta: Caution should be exercised setting the
tube current (output power) of the Argon/2 laser, (blue 488 nm).
Maximum laser tube current on the Argon/2 (488 nm laser) has become
un-calibrated and can be set above 8.1 Amps. This level should not be
exceeded. To this end the output level should be gently ramped up to
the desired power rather than rapidly dragging the slider to high output
levels. Rapid dragging of the power slider tends to set the tube current to
levels about 8.1 Amps. This laser is showing signs that it is going to
fail. Replacement had been scheduled for 8:00 to 10:00 AM on April 13.
However Fedex declined delivery of the unit, so installation could not
- Feb 18 - Leica SP2 UV laser: A loaned UV
laser head has been installed and aligned. Dapi, Alexa 350, Fura-2,
Cascade Blue, etc. may be imaged again.
- Also note that the "blue" Argon Laser (488 nm) must never be turned off
without first turning the start key to off, and letting the laser cool down
before turning the cooling fan and power supply (red laser switch) off.
Check the cooling exhaust temperature (upper temperature gage on the
thermometer on the conduit) which will be less than 5oC above the
ambient temperature (lower gage on the thermometer).
- Feb 16 - Leica SP2 UV laser: The
EnterpriseII UV laser head is scheduled to ship back from Coherent on April
1. Once the laser is received in Chapel Hill it will be refitted back
into the confocal, and the confocal is expected to be unavailable for about
half a day
- Jan 22 - Leica SP2 UV laser: The
EnterpriseII UV laser has finally been received by Coherent of Santa Clara,
CA as of Jan 19. Expected repair turn around time is 2 months.
Status of the laser repair will be posted here as it becomes available.
- Dec 31 - Leica SP2 UV laser: The UV
laser has been shipped to Coherent for repair. UV scanning will be
unavailable until the unit is shipped back. Expected turn around time is
2 months. Scanning in the visible wavelengths is still available.
Please power down the green 561 nm laser by pushing the red button on the
power supply unit when finished scanning.
- Dec 10 - Leica SP2 UV laser: Coherent
turned out not to be available to conduct the planned alignment on the
EnterpriseII laser. This will be attempted on Dec 13. Coherent
California/Tennessee 800-367-7890 case# 349 366
- Nov 23 - LeicaSP2 confocal: The UV laser
can only operate at very low output power ( <3 mW) due to over heating. Service
is being arranged. It can still be used with appropriate power adjustments.
1. Turn power knob on the control box
1/4 to 1/2 turn; be sure the "Current Limit" light is not illuminated.
2. Adjust the 364nm AOTF power as
high as necessary (up to 100%) for a good image at Gain setting of 500-650
- Nov 1 - Leica Inverted Microscope: The polarizer has been replaced.
Users should remember to have the analyzer out of the light path when doing
fluorescence. The analyzer will cut fluorescence by 50%. (It
also is an expensive neutral density filter and is best not be used for that
- Oct 18 - LeicaSP2 confocal: Laser and scanhead fibers realigned.
AOTF retuned. Output of all visible laser lines have been
substantially increased at the stage.
- Oct 13 -
Leica Inverted standard microscope: The polarizer material of the
analyzer for DIC is wrinkling. This is reducing the quality of DIC (Nomarski)
images slightly. Only DIC imaging is affected, and since standard
transmitted light and fluorescence do not use the analyzer these modes of
imaging are not affected. A new analyser has been ordered.
Delivery is expected in 2 to 3 weeks.
- Sep 30 - Booking Calendars: It may not be
possible to display or make bookings when using Internet Explorer under
WindowsXP. Try 1. reducing the security level for Internet Explorer
(done in the control panel), 2. use another browser such as Netscape/Mozilla,
3. book from another computer which does not use WindowsXP, 4. call the
facility or e-mail to make a booking. Once the computer person gets a computer running WindowsXP, he
can look into diagnosing and fixing this problem. In the meantime please follow
one of the workarounds described above.
- Sep 29 - LeicaSP2 confocal: The AOBS was realigned at 561 nm.
Reflections should be reduced when scanning TexasRed, Rhodamine, Alexa 594,
- Sep 28 - Leica SP2 confocal: Condenser is
fixed. Swinging pin is reseated and glued in place.
- Sep 27 - Leica SP2 Confocal: Currently it
is not possible to align the transmitted light condenser due to minor damage.
The problem has no impact on confocal or widefield florescence viewing or
scanning. However it is not possible to perform transmitted light
imaging during confocal scanning or center the condenser for DIC imaging.
The unit should be operational again on Sep 28.
- Sep 24 - Zeiss 510 confocal: the x-axis galvanometer (horizontal
scanning driver) has been replaced and the machine recalibrated.
- Sep 20 - Zeiss 510 confocal: the x-galvo is
giving some minor problems. Images may jitter in the x direction.
The x-galvanometer motor will be repaired on September 24.
- Sep 2 - Leica SP2: The 561 nm green solid state laser has
been installed. All laser lines are working. Note that the
new green 561 nm laser has a simple on/off buttons. No cooling is
required. The 561 AOTF should be run at roughly 5% (cf. 50% to 100% for
the old Krypton 568 nm laser)
- May 12 - Zeiss 510 Meta: The channel 2
emission band pass filter 475-505 nm information has been entered into the
confocal database and is displayed correctly in the filter setup window in the
LSM 510 software (and on
the web page). This filter is intended for CFP detection.
Warning listed at April 28 no longer applies.
- April 28: Zeiss 510 Meta: The channel 2 emission filter (505-550
nm) has been replaced with a 475 to 505 nm emission filter (Chroma HD490/30m).
CH2 (PMT 2) filter display still shows 505-550 nm
when actually the 475 to 505 nm filter is in place. This needs to be
updated. The purpose of having a 475 to 505 emission filter is to help
with simultaneous CFP & YFP detection using PMTs (not the meta) the 458 and
514 nm laser lines for excitation. Prior to this upgrade the closest
filter for CFP imaging was the 475 to 525 nm emission filter. This
filter is undesirable in that it lets reflection from the 514 laser line
through to the PMT detector. For more information contact Michael
- Zeiss 510 confocal: The Axioscope will be
sent out for service due to corrosion on the turret mirror. This mirror
requires replacement, which can only be done at the factory since it must be
critically aligned. Confocal will be unavailable for 7-9 weeks at a date
yet to be decided.